OpenStudy (anonymous):

if logbase(6)4=.774 and logbase(6)10=1.285 use the properties of logs to approximate the value of logbase(6)40?

OpenStudy (ipwnbunnies):

This property works with any base, but they have to be the same base: \[\log (a) + \log(b) = \log(a*b)\]